# Good News for Polynomial Root-finding

### v35

@@ B.2 Lower bounds on the bit operation complexity
Observation B.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$. Sum these bounds for $j=1,\dots,d$ and obtain
Corollary B.2.
One must access at least $0.5(d-1)db$ bits of the coefficients of $p$ and perform at least $0.25(d-1)db$ bit operations to approximate even a single zero of $p$ within $2^{-b}$.
Likewise, deduce lower bounds on the number of bit operations required for Problem 0.
Corollary B.3.
To solve Problem 0 and even to approximate within $\epsilon$ a single zero of $p$ in the disc $D$, one must access at least $0.5b\max\{d+1,(m-1)m\}$ bits of input coefficients and perform at least $0.25b\max\{d+1,(m-1)m\}$ bit operations for $b=\log_{2}(R/\epsilon)$ and $R=|c|+\rho$.
The lower bounds of this subsection imply that the algorithm of [92, 99] and its extension supporting Thm. B.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylog factors or, in simple terms.
B.3 Linking Factorization and Root-finding (with more details)
Next we prove Thm. B.2, by applying the results of [130, Cor. 4.3 and Sec. 19].
@@ B.3 Linking Factorization and Root-finding (with more details)
$\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d-1)db$ (B.2)
of [92, 99] and the following result (cf. [130, Thm. 19.1]).
Theorem B.4.
@@
Based on [130, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
Furthermore, in view of (B.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.4), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
[132, Thm. 2.7], which we reproduce in Sec. C.4 as Thm. C.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
Schönhage in [130, Sec. 19] has greatly strengthened the estimate of Thm. B.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [130, Eqn. (19.8)]) the following result.
Lemma B.1.
Assume that the polynomial $p(x)$ is monic, the values $u_{j}=1$ for all $j$ and complex $v_{1},\dots,v_{d}$ satisfy bound (B.1), and $|z_{j}-v_{j}|\leq r:=2^{2-b/d}$ for all $j$. Furthermore, write $\Delta_{j,1}:=|v_{j}-v_{1}|$ for $j=2,3,\dots,d$ and suppose that $1.25\Delta_{j,1}\leq r$ for $j=2,3,\dots,d$ and $\Delta_{1}:=\min_{j}\Delta_{j,1}>r$. Then $|z_{1}-v_{1}|\leq 2^{2-b/m}/|\Delta_{1}-r|^{\frac{d-m}{m}}$.
If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma B.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
By extending these estimates we strengthen Thm. B.4 as follows.
Corollary B.4.
Under bounding cluster sizes of Def. B.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$. Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{1}(b,d)\leq\mathbb{B}_{2}(\tilde{O}(b+d),d)$.
For converse reduction of Problem F to Problem 0 recall [130, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
Theorem B.5.
@@
Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$. For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the zero $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F. Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm. B.5. Hence the approximation errors of all the $d$ zeros of $p$ together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
Corollary B.5.
@@
It follows that the complexity estimate of Thm. 1.3 can be also applied to the solution of Problem F of factorization provided that $b=O(b-d-\log(d))$. Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf. Observation B.1).
B.4 Two root-finders of 2016 and 2018
@@
Implementation of the root-finder of [22] by Imbach et al in [61] provided no empirical support to optimality claimed by Becker et al because the root-finder of [60], using some novel techniques from [108] (see [60, Sec. 1.2]), runs much faster. Unfortunately, formal support for optimality claimed by Becker et al has been missing as well.
@@
[…]
Theorem B.6.
@@
To apply Pellet’s theorem to $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$. Hence, by virtue of Observation B.2, one must process at least $0.5(d-1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
@@
[…]
Appendix C Recursive factorization of a polynomial and extensions to root-finding and root isolation
@@
C.1 Auxiliary norm bounds
@@
The leftmost bound was proved by Mignotte in [75]. The rightmost bound follows from Lemma C.2. ∎
Remark C.1.
[132, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma C.3. From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta:=\zeta_{q}:=\exp(\frac{2\pi{\bf i}}{q})$ denoting a primitive $q$-th root of unity, one can see some limitations on strengthening this bound further.
C.2 Errors and complexity
@@
$p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (C.1)
Next, by following [130, Sec. 5], we estimate the norm of the residual polynomial
$\Delta^{\ast}=p^{\ast}-p.$ (C.2)
@@
$\Delta_{k}=|p-f_{1}\cdots f_{k}|\leq\frac{k\epsilon|p|}{d},$ (C.3)
$\Delta=|f_{1}-fg|\leq\epsilon_{k}|f_{1}|,$ (C.4)
@@
$\epsilon_{k}d2^{d}\prod_{j=1}^{k}|f_{j}|\leq\epsilon.$ (C.5)
@@
$|\Delta_{k+1}|=|p-fgf_{2}\cdots f_{k}|\leq\frac{(k+1)\epsilon|p|}{d}.$ (C.6)
@@
$\Delta_{k+1}=|p-f_{1}\cdots f_{k}+(f_{1}-fg)f_{2}\cdots f_{k}|.$ Apply Lemma C.1 and deduce that $\Delta_{k+1}\leq\Delta_{k}+|(f_{1}-fg)f_{2}\cdots f_{k}|$ and, furthermore, that
@@
Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (C.3)-(C.5) and Lemmas C.1 and C.3 and obtain (C.6). ∎
Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (C.6) turns into (C.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (C.4), apply Thm. C.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (C.1) with
@@
for $\Delta^{*}$ of (C.2). Let us call this computation Recursive Splitting Process provided that it begins with $k=1$ and $f_{1}=p$ and ends with $k=d$.
Theorem C.2.
In order to support (C.3) for $0<\epsilon\leq 1$ and $j=1,2,\ldots,d$ in the Recursive Splitting Process it is sufficient to choose $\epsilon_{k}$ in (C.4) satisfying
$\epsilon_{k}~{}d~{}2^{2d+1}\leq\epsilon~{}\textrm{for~{}all}~{}k.$ (C.7)
@@
Prove bound (C.3) by induction on $j$. Clearly, the bound holds for $k=1$. It remains to deduce (C.6) from (C.3) and (C.7) for any $k$. By first applying Lemma C.3 and then bound (C.3), obtain
@@
The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (C.7) ensures (C.5), and then (C.6) follows by virtue of Thm. C.1.
Remark C.2.
@@
C.3 Overall complexity of recursive factorization and extension to Root Isolation
The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors. The algorithm of [99] adopts recursive factorization of [130] but decreases the overall complexity bound by roughly a factor of $d$. This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11. Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [99], and so the overall complexity of root-finding is proportional to that of the first splitting. In contrast, the degrees of a factor in any splittin in [130] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
C.4 From factors to roots
Theorem C.3.
[132, Thm. 2.7]. Suppose that
@@ C.4 From factors to roots
By virtue of Thm. C.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec. C.3 by a factor of $d$.
Corollary C.1.
The bit operation complexity of the solution of Problem 0. Given a polynomial $p$ of degree $d$ and a real $b^{\prime}\geq 7d$, one can approximate all zeros of that polynomial within the error bound $2^{-b^{\prime}}$ by using $O(\mu(b^{\prime}+d)d^{2}\log(d)(\log^{2}d+\log(b^{\prime})))=\tilde{O}((b^{\prime}+d)d^{2})$ bit operations.
Remark C.3.
The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$. If its roots are isolated from each other, then the upper estimates of Thm. C.3 and Cor. C.2 can be decreased by a factor of $d$ (cf. [99]).
C.5 Polynomial Root Isolation
Based on his results recalled in this appendix, Schönhage in [130, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
@@ C.5 Polynomial Root Isolation
Remark C.4.
Our Problem $*{}^{*}$ can be viewed as the Problem of Polynomial Root Cluster Isolation. It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem $*{}^{*}$ as well as the Algorithms of [92, 99] solve Problem ISOL within a record complexity bounds.
In [130, Sec. 20] Schönhage reduces to factorization Problem F the root-isolation problem of computing $d$ disjoint discs each covering a single simple root of a square-free polynomial with integer coefficients. Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [74, 75, 76]), and Schönhage has estimated that solution of Problem F (e.g., by means of the algorithms of [130, 90, 99]) implies isolation provided that
$b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (C.8)
@@
Appendix D Newton’s, Weierstrass’s, and Ehrlich’s iterations
D.1 Contents
In the next subsection we comment on convergence of the three listed iterative root-finders and on their stopping criterion. In Sec. D.3 we discuss the policy of doubling precision, the resulting dynamic partition of the roots into tame and wild, and the deflation of a wild factor. In Sec. D.4 we accelerate these iterations by incorporating FMM. In Sec. D.5 we compute a disc that contains all zeros of a black box polynomial $p$, which one needs in customary initialization of root-finding for a black box polynomial by means of Newton’s iterations (cf. [138]). Customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations (see Sec. D.6).
D.2 Definition
1. Definition. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [68, 55, 126, 138, 143, 29, 64, 36, 3, 21]):
$z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (D.1)
$N(x)=W(x)=E_{j}(x)=0~{}{\rm for~{}all}~{}j~{}{\rm if}~{}p(x)=0$ (D.2)

### v36

@@ 24.2 Lower bounds on the bit operation complexity
Observation 24.2 implies that one must access the coefficient $p_{d-j}$ within $2^{-jb}$ to approximate within $2^{-b}$ a $j$-multiple zero of $p$ or its zero in a cluster of $j$ zeros, and so a root-finder must process at least $jb$ bits of the coefficient $p_{d-j}$. Sum these bounds for $j=1,\dots,d$ and obtain
Corollary 24.2.
One must access at least $0.5(d+1)db$ bits of the coefficients of a monic polynomial $p$ and perform at least $0.25(d+1)db$ bit operations to approximate even its single within $2^{-b}$.
The lower bounds of this subsection imply that the algorithm of [96, 103] and its extension supporting Thm. 24.1 solve Problems 0 and F of large size by using optimal number of bit operations up to polylogarithmic factors.
24.3 Linking Factorization and Root-finding (with more details)
Next we prove Thm. 24.2, by applying the results of [134, Cor. 4.3 and Sec. 19].
@@ 24.3 Linking Factorization and Root-finding (with more details)
$\mathbb{B}_{0}(b,d)\geq 0.5(d+1)b~{}{\rm and}~{}\mathbb{B}_{F}(b,d)\geq 0.25(d+1)db$ (24.2)
of [96, 103] and the following result (cf. [134, Thm. 19.1]).
Theorem 24.4.
@@
Based on [134, Cor. 4.3] we can also readily relax the assumption that $p(x)$ is monic.
Furthermore, in view of (24.2), relaxing the bound $r_{1}\leq 1$ would little affect the latter estimate because we can apply map (1.3), which changes the error bound of Problem 0 by at most $\log_{2}(\max\{\rho,1/\rho\})$, for $\rho\geq 1/2^{b}$.
[136, Thm. 2.7], which we reproduce in Sec. 25.4 as Thm. 25.3, shows an alternative way to relaxing both assumptions – that $r_{1}\leq 1$ and $p(x)$ is monic.
Schönhage in [134, Sec. 19] has greatly strengthened the estimate of Thm. 24.4 unless “the zeros of $p$ are clustered too much". Namely, he proved (see [134, Eqn. (19.8)]) the following result.
Lemma 24.1.
Assume that the polynomial $p(x)$ is monic, the values $u_{j}=1$ for all $j$ and complex $v_{1},\dots,v_{d}$ satisfy bound (24.1), and $|z_{j}-v_{j}|\leq r:=2^{2-b/d}$ for all $j$. Furthermore, write $\Delta_{j,1}:=|v_{j}-v_{1}|$ for $j=2,3,\dots,d$ and suppose that $1.25\Delta_{j,1}\leq r$ for $j=2,3,\dots,d$ and $\Delta_{1}:=\min_{j}\Delta_{j,1}>r$. Then $|z_{1}-v_{1}|\leq 2^{2-b/m}/|\Delta_{1}-r|^{\frac{d-m}{m}}$.
If $1/|\Delta-r|^{\frac{d-m}{m}}=2^{\tilde{O}(d)}$, then Lemma 24.1 implies that $|z_{1}-v_{1}|\leq 1/2^{b/m-\tilde{O}(d)}$, and we can similarly estimate $|v_{j}-z_{j}|$ for all $j=1,2,\dots,d$.
By extending these estimates we strengthen Thm. 24.4 as follows.
Corollary 24.3.
Under bounding cluster sizes of Def. 24.1, it holds that $|z_{j}-v_{j}|\leq 1/2^{b/m-\tilde{O}(d)},~{}j=1,2,\dots,d$. Hence if $m=O(1)$ is a constant, then a solution of Problem 0 within an error bound $1/2^{b}$ can be recovered at a dominated bit operation cost from the solution of Problem F with a precision $1/2^{\bar{b}}$ for a sufficiently large $\bar{b}=\tilde{O}(b+d)$, and so $\mathbb{B}_{1}(b,d)\leq\mathbb{B}_{2}(\tilde{O}(b+d),d)$.
For converse reduction of Problem F to Problem 0 recall [134, Cor. 4.3] in the case of $k=2$, covering the product of two polynomials:
Theorem 24.5.
@@
Now consider the products $p=gh$ for $g=x-z_{j}$, for $j=1,\dots,d$. For a fixed $j$, an error of at most $\epsilon/2^{d}$ for the root $z_{j}$ contributes at most $\epsilon||h||_{1}$ to overall factorization error of Problem F. Therefore, $||h||_{1}\leq 2^{d-1}||p||_{1}/||x-z_{j}||_{1}\leq 2^{d-1}||p||_{1}$ by virtue of Thm. 24.5. Hence the approximation errors of all the $d$ roots together contribute at most $2^{d-1}d\epsilon$ to the relative error of the solution of Problem F (ignoring dominated impact of higher order terms), and we arrive at
Corollary 24.4.
@@
It follows that the complexity estimate of Thm. 1.3 can be also applied to the solution of Problem F of factorization provided that $b=O(b-d-\log(d))$. Then again, the bit operation complexity of this solution of Problem F is optimal (up to polylogarithmic factors) under the random root model provided that $d=\tilde{O}(b)$ (cf. Observation 24.1).
24.4 Two root-finders of 2016 and 2018
@@
Implementation of the root-finder of [23] by Imbach et al in [64] provided no empirical support to optimality claimed by Becker et al because the root-finder of [63], using some novel techniques from [112] (see [63, Sec. 1.2]), runs much faster. Unfortunately, formal support for optimality claimed by Becker et al has been missing as well.
@@
Can one salvage the claim by Becker et al in another way? Next we argue that this can be problematic because the algorithms of [23, 24] for Problem 1m are some variants of subdivision iterations. The basic tool and the main algorithmic novelty of the variant of subdivision in [96, 103] are e/i tests based on the following theorem by Pellet [88], which extended the previous results by Cauchy and Rouché:
Theorem 24.6.
@@
To apply Pellet’s theorem to $t(x)$ one must compute the values $|\frac{t_{i}}{t_{d}}|$ for $i=0,1,\dots,d-1$. Hence, by virtue of Observation 24.2, one must process at least $0.5(d+1)db$ bits of the coefficients of $t(x)$ to support approximation of its zeros within $1/2^{b}$.
@@
Already a single subdivision iteration can perform more than $4m$ soft e/i tests; this means that one must process more than $2(d+1)dmb$ bits of memory and $(d+1)dmb$ bit-operations to solve Problem 1m by applying such a Pellet-based root-finder.
25 Recursive Factorization; extensions to Root-finding and Root Isolation
@@
25.1 Auxiliary norm bounds
@@
The leftmost bound was proved by Mignotte in [78]. The rightmost bound follows from Lemma 25.2. ∎
Remark 25.1.
[136, Lemma 2.6] states with no proof a little stronger bound $\prod_{j=1}^{k}|f_{j}|\leq 2^{d-1}|p|$ under the assumptions of Lemma 25.3. From various factors of the polynomial $p(x)=x^{d}-1$ such as $\prod_{j=1}^{\frac{d}{2}}(x-\zeta_{d}^{j})$ for even $d$ and $\zeta:=\zeta_{q}:=\exp(\frac{2\pi{\bf i}}{q})$ denoting a primitive $q$-th root of unity, one can see some limitations on strengthening this bound further.
25.2 Errors and complexity
@@
$p\approx p^{\ast}=p^{\ast}(x)=\prod_{j=1}^{d}(u_{j}x+v_{j}).$ (25.1)
Next, by following [134, Sec. 5], we estimate the norm of the residual polynomial
$\Delta^{\ast}=p^{\ast}-p.$ (25.2)
@@
$\Delta_{k}=|p-f_{1}\cdots f_{k}|\leq\frac{k\epsilon|p|}{d},$ (25.3)
$\Delta=|f_{1}-fg|\leq\epsilon_{k}|f_{1}|,$ (25.4)
@@
$\epsilon_{k}d2^{d}\prod_{j=1}^{k}|f_{j}|\leq\epsilon.$ (25.5)
@@
$|\Delta_{k+1}|=|p-fgf_{2}\cdots f_{k}|\leq\frac{(k+1)\epsilon|p|}{d}.$ (25.6)
@@
$\Delta_{k+1}=|p-f_{1}\cdots f_{k}+(f_{1}-fg)f_{2}\cdots f_{k}|.$ Apply Lemma 25.1 and deduce that $\Delta_{k+1}\leq\Delta_{k}+|(f_{1}-fg)f_{2}\cdots f_{k}|$ and, furthermore, that
@@
Combine the latter inequalities and obtain $\Delta_{k+1}\leq\Delta_{k}+\Delta|f_{2}\cdots f_{k}|$. Combine this bound with (25.3)-(25.5) and Lemmas 25.1 and 25.3 and obtain (25.6). ∎
Write $f_{1}:=f$ and $f_{k+1}=g.$ Then (25.6) turns into (25.3) for $k$ replaced by $k+1.$ Now compute one of the factors $f_{j}$ as in (25.4), apply Thm. 25.1, then recursively continue splitting the polynomial $p$ into factors of smaller degrees, and finally arrive at factorization (25.1) with
@@
for $\Delta^{*}$ of (25.2). Let us call this computation Recursive Splitting Process provided that it begins with $k=1$ and $f_{1}=p$ and ends with $k=d$.
Theorem 25.2.
In order to support (25.3) for $0<\epsilon\leq 1$ and $j=1,2,\ldots,d$ in the Recursive Splitting Process it is sufficient to choose $\epsilon_{k}$ in (25.4) satisfying
$\epsilon_{k}~{}d~{}2^{2d+1}\leq\epsilon~{}\textrm{for~{}all}~{}k.$ (25.7)
@@
Prove bound (25.3) by induction on $j$. Clearly, the bound holds for $k=1$. It remains to deduce (25.6) from (25.3) and (25.7) for any $k$. By first applying Lemma 25.3 and then bound (25.3), obtain
@@
The latter bound cannot exceed $2^{d+1}|{p}|$ for $k\leq d$, $\epsilon\leq 1$. Consequently (25.7) ensures (25.5), and then (25.6) follows by virtue of Thm. 25.1.
Remark 25.2.
@@
25.3 Overall complexity of Recursive Factorization and extension to Root Isolation
The overall complexity of recursive factorization is dominated by the sum of the bounds on the complexity of all splittings into pairs of factors. The algorithm of [103] adopts recursive factorization of [134] but decreases the overall complexity bound by roughly a factor of $d$. This is due to additional techniques, which ensure that in every recursive splitting of a polynomial into the product of two factors the ratio of their degrees is at most 11. Consequently, $O(\log(d)$ recursive splitting steps are sufficient in [103], and so the overall complexity of root-finding is proportional to that of the first splitting. In contrast, the degrees of a factor in any splitting
in [134] can have low, e.g., 1, and so there can be allows up to $d-1$ splittings overall, with each of the first $0.5d$ of them almost as expensive as the very first one.
25.4 From factors to roots
Theorem 25.3.
[136, Thm. 2.7]. Suppose that
@@ 25.4 From factors to roots
By virtue of Thm. 25.3 for $b^{\prime}=O(bd)$ we can bound the bit operation complexity of the solution of Problem 0 by increasing the estimate for the complexity of factorization in Sec. 25.3 by a factor of $d$.
Corollary 25.1.
The bit operation complexity of the solution of Problem 0. Given a polynomial $p$ of degree $d$ and a real $b^{\prime}\geq 7d$, one can approximate all its $d$ zeros within the error bound $2^{-b^{\prime}}$ by using $O(\mu(b^{\prime}+d)d^{2}\log(d)(\log^{2}d+\log(b^{\prime})))=\tilde{O}((b^{\prime}+d)d^{2})$ bit operations.
Remark 25.3.
The study in this section covers the solution of Problem 1 for the worst case input polynomial $p$. If its roots are isolated from each other, then the upper estimates of Thm. 25.3 and Cor. 25.2 can be decreased by a factor of $d$ (cf. [103]).
25.5 Polynomial Root Isolation
Based on his results recalled in this section, Schönhage in [134, Sec. 20] estimates the bit operation complexity of the following well-known problem by reducing it to Problem F.
@@ 25.5 Polynomial Root Isolation
Remark 25.4.
Our Problem 0 can be viewed as the Problem of Polynomial Root Cluster Isolation. It turns into Problem ISOL if all the coefficients of $p$ are integers and the roots are simple, and so our algorithms for Problem 0 as well as the Algorithms of [96, 103] solve Problem ISOL within a record complexity bounds.
In [134, Sec. 20] Schönhage reduces to factorization Problem F the root-isolation problem of computing $d$ disjoint discs each covering a single simple root of a square-free polynomial with integer coefficients. Problem ISOL is readily solved if the zeros are approximated within $\epsilon=1/2^{b}<0.5\delta$ for $\delta$ denoting the minimal pairwise distance between the zeros of $p$ (see the known lower bounds on $\delta$ in Kurt Mahler and Maurice Mignotte [77, 78, 79]), and Schönhage has estimated that solution of Problem F (e.g., by means of the algorithms of [134, 94, 103]) implies isolation provided that
$b=O((\tau+\log(d))+d)~{}{\rm and}~{}\tau=\log_{2}\max_{i=0}^{d}|p_{i}|.$ (25.8)
@@
26 Newton’s, Weierstrass’s, and Ehrlich’s iterations
26.1 Contents
In the next subsection we recall Newton’s, Weierstrass’s, and Ehrlich’s iterations, comment on their convergence, and acceleration by means of applying fast multipoint evaluation, which is dramatic but involves the coefficients of $p$. In Sec. 26.3 we cover mre limited acceleration of these iterations by means of incorporating FMM, which can be applied to a black box polynomial. In Secs. 26.5 and 27) we cover initialization of Newton’s, Weierstrass’s, and Ehrlich’s iterations, required for their fast global empirical convergence. For Newton’s iterations we only need to compute a disc that contains all zeros of a polynomial $p$ (cf. [71, 58, 143, 130, 141]). This is given by the well-known estimates via the coefficients, but we present simple but novel recipes in the case of black box polynomial $p$. The customary initialization is more demanding for Weierstrass’s and Ehrlich’s iterations but is still fast for polynomials represented with their coefficients, as we recall in Sec. 27), where we also point out how this initialization can be extended to fast real and nearly real root-finding. It is still a challenge to extend such an initialization to the case of a black box polynomial.
26.2 Definitions
1. Definitions. Consider Newton’s, Weierstrass’s (aka Durand-Kerner’s), and Ehrlich’s (aka Aberth’s) iterations [71, 58, 130, 143, 148, 31, 67, 38, 3, 22]):
$z\leftarrow z-N(z),~{}z_{j}\leftarrow z_{j}-W(z_{j}),~{}z_{j}\leftarrow z_{j}-E_{j}(z_{j}),~{}j=1,\dots,d,~{}{\rm where}$ (26.1)
$N(x)=W(x)=E_{j}(x)=0~{}{\rm for~{}all}~{}j~{}{\rm if}~{}p(x)=0$ (26.2)